OpenStudy (anonymous):

A sports nutritionist recommends that exercisers drink 2 cups of water or other liquid before a race or workout and 1 cup for every 15 minutes they exercise during a workout to avoid dehydration. a) Suppose a runner takes 3 hours and 15 minutes to complete a marathon. How many cups of fluid should the runner consume for this event? b) Let F represent cups of fluid and N represent number of 15-minute intervals spent exercising, and write an equation that can be used to find the recommended number of cups of fluid to drink while exercising. c) sketch a graph of the equation

OpenStudy (anonymous):

some please help me !

OpenStudy (anonymous):

a) 14 cups if my working out is correct.

OpenStudy (anonymous):

1hour = 4x15min so 3hours = 3x4x15min=12 cups and then u had another 15 min to complete the race + the 2 cups before the race which gives us: 12+1+2= 15 cups of water

OpenStudy (anonymous):

F=Cups N=every 15minutes interval M=the starting Cups So, F=N+M So if i run for an hour, 60min, which is 4 N's i will drink: F=4+2 -> 6 cups

OpenStudy (anonymous):

thankyou !